The measure of one angle of a triangle is 26 degrees less than twice the measure of another angle. the measure of the third angle is 35. find the measure of the other two angles
Answer by rm29924(97)

The sum of angles in a triangle = 180 degrees
one angle is 35
let x=another angle
the third angle is 2x-26
35+x+2x-26=180
combine like terms
9+3x=180
subtract 9 from both sides
3x=171
divide both sides by 3 to get the value of x
x=57
2(57)-26=88
the other two angles are 57 degrees and 88 degrees
check work
35+57+2(57)-26=180
35+57+114-26=180
180=180